# A multi-band study of pulsar glitches with Fermi-LAT and Parkes

### v1

@@ Abstract
Pulsar glitch is a phenomenon characterized by abrupt changes in the spin period over less than a minute. We present a comprehensive analysis of glitches in four gamma-ray pulsars by combining the timing observation data of Fermi Large Area Telescope (Fermi-LAT) and Parkes 64 m radio telescope. The timing data of five pulsars, namely PSRs J1028--5819, J1420--6048, J1509--5850, J1709--4429 (B1706--44) and J1718--3825, spanning over 14 years of observations for each, are examined. A total of 12 glitches are identified in four pulsars, including a previously unreported glitch. That is, a new small glitch is identified for PSR J1718--3825 in MJD $\sim$ 59121(8), and the fractional glitch size was $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$. For PSR J1420--6048, our investigation confirms the existence of two linear recovery terms during the evolution of $\dot{\nu}$ subsequent to glitches 4, 6 and 8, and identified an exponential recovery process in glitch 8, with $Q=0.0131(5)$, $\tau_{\rm d}=100(6)$ d. Regarding the fourth glitch of PSR J1709--4429, our analysis reveals the presence of two exponential recovery terms with healing parameters and decay time-scales $Q$1=0.0104(5), $\tau_{\rm d1}=72(4)$ d and $Q$2 = 0.006(1), $\tau_{\rm d2}=4.2(6)$ d, respectively. For the remaining previously reported glitches, we refine the glitch epochs and glitch observables through precise fitting of the timing residual data. We extensively discuss how multi-band data of glitches can help better characterize the glitch recoveries and constrain the underlying physics of glitch events. We demonstrate that the accumulation of observational data reveals the rich complexity of the glitch phenomenon, aiding in the search for a well-established interpretation.
@@ 1 Introduction
PSR RA DEC $P$ $\dot{P}$ $B_{\rm s}$ Age DM Fermi-LAT span Parkes span
@@
Gamma-ray pulsars  [^1] Except millisecond pulsars, characterized as youthful neutron stars with high spin-down rates active in the gamma-ray band of emitted radiation, are good target sources for searching and studying glitches (Sokolova & Panin, 2022). More than 150 young gamma-ray pulsars have been reported so far, with more than 80 among them having exhibited radio emissions (Smith et al., 2023). Studying glitches in gamma-ray pulsars holds the potential to not only glean valuable insights into the internal structure of neutron stars but also to offer a window into the dynamic evolution of the magnetosphere (Kou et al., 2018).
Currently, there has been some progress in research focused on glitches in gamma-ray pulsars. PSR J2021+4026 is the only glitching pulsar for which a gamma-ray flux jump has been observed (Allafort et al., 2013; Takata et al., 2020). This phenomenon has been linked to glitches, with studies by Ng et al. (2016); Zhao et al. (2017) indicating a plausible association between gamma-ray flux jumps and substantial shifts in the pulse profile of PSR J2021+4026. In a similar manner, Lin et al. (2021) reported the detection of a new glitch in PSR J1420--6048 using Fermi-LAT 11 yr timing data and found that the gamma-ray pulse profile changed significantly within certain specific time ranges. Gügercinoğlu et al. (2022) identified 20 glitches in 4 gamma-ray pulsars, and made a relatively accurate prediction of the waiting time of the next glitch through the vortex creep model.
In this paper, we study the glitches of five gamma-ray pulsars by combining the timing observation data of the Fermi-LAT and Parkes telescopes in the past 14 yr. We have identified 12 glitches in total and detected a new small glitch at PSR J1718--3825. Besides, we update the epoch and other parameters of the previously reported 11 glitches.
The structure of this paper is as follows: In Section 2, we briefly introduce the timing observations of Fermi-LAT and Parkes 64 m radio telescope and give an overview of the glitch data processing procedure. One superfluid model, namely the vortex creep model, that is used in glitch fitting is introduced in Section 3, following which we relate the glitch observables to the dynamical and structural properties of the neutron star. The results and underlying physical process of pulsar glitches studied in the paper are detailed in Section 4. We then discuss our results more extensively in Section 5 and finally summarize our findings in Section 6.
@@ 2 Observation and analysis
In this work, we analysed the timing observation data of Fermi-LAT and Parkes 64 m radio telescope in Australia from 2007 to 2022 for PSRs J1028--5819, J1420--6048, J1509--5850, J1709--4429 and J1718--3825. Table 1 presents the detailed parameters of pulsars. The first three columns are the pulsar name (J2000) and the position [right ascension (RA) and declination (DEC)] in the celestial coordinate system. The fourth and fifth columns are the pulsar spin period ($P$) and its first time derivative ($\dot{P}$). The sixth and seventh columns are the inferred surface dipole magnetic field strength at the stellar equator ($B_{\rm s}=3.2\times 10^{19}\sqrt{P\dot{P}}$ G) and the characteristic age ($\tau_{\rm c}=P/2\dot{P}$) of pulsars. The last three columns are the pulsar’s dispersion measure (DM) and the timing observation data time span of the Fermi-LAT and the Parkes telescopes for the corresponding sources.
@@
We also collected the timing observation data of these five gamma-ray pulsars from 2007 to 2020 on the Parkes pulsar data archive [^3] https://data.csiro.au/domain/atnf. Manchester et al. (2013) has introduced the timing observation of Parkes 64 m radio telescope in detail. Briefly, Parkes observation was carried out at 20 cm band, mainly using the multi-beam receiver, with a centre frequency of 1369 MHz and bandwidth of 256 MHz (Staveley-Smith et al., 1996; Hobbs et al., 2020). The Parkes Digital Filter Bank system used includes PDFB1, PDFB2, PDFB3 and PDFB4, and the generated file has 1024 frequency channels, each channel bandwidth is 0.25 MHz. The observation interval of each pulsar is 2–4 weeks, the observed sub-integration time was 30 s, and the integration time was 2–15 mins.
@@
For radio timing data, PSRCHIVE was used to eliminate the dispersion effects and radio frequency interference (RFI), while superimposing to obtain an average integrated pulse profile for each observation (Hotan et al., 2004; van Straten et al., 2012). The timing data were partitioned based on glitch epoch, and then PSRADD was used to superimpose the average pulse profile of each observation in the polarization, time domain and frequency to obtain the corresponding total integrated pulse profile. The total integrated pulse profile with the highest signal-to-noise ratio (S/N) was selected as the standard pulse profile. The PAT tool within PSRCHIVE was then used to cross-correlate each mean profile with the standard profile, obtaining highly accurate pulse times of arrival (ToAs; Taylor 1992). Finally, the Jet Propulsion Laboratory planetary ephemeris DE440 (Park et al., 2021) and the Barycentric Coordinate Time (TCB) were used to convert these local ToAs to SSB, thereby eliminating the effect of the Earth motion on the ToAs.
@@
We combined the gamma-ray and radio ToAs, and used the high-precision timing analysis software packages TEMPO2 (Hobbs et al., 2006; Edwards et al., 2006) to achieve the timing solution. We introduced two parameters, EFAC and EQUAD, to correct the uncertainty of each ToA acquired from different telescopes and terminals (Wang et al., 2015). Among them, EFAC is a free parameter that adjusts unaccounted instrumental errors, and EQUAD accounts for additional time-independent white noise processes (Lentati et al., 2014). The values of these two parameters were determined by the EFACEQUAD plug-in of TEMPO2, and this method did not involve Bayesian techniques. After correction, the relationship between the initial uncertainty ($\delta$) of ToAs and its new value (${\delta}_{\rm s}$) was defined as ${\delta}_{\rm s}^{\rm 2}={\rm EFAC}^{2}\times(\delta^{2}+{\rm EQUAD}^{2})$. After adding the error correction of ToA, the reduced chi-square values of each fitting region are distributed between 0.95–1.05.
The pulsar position is one important factor that affects the precision of glitch parameter, but for some sources, the position obtained from PSRCAT is not accurate enough. For these pulsars, we used the "Cholesky" method to redefine the position of the pulsars (Coles et al., 2011). To do this, we selected a sufficiently long length of data that not include any glitch, invoked the TEMPO2 plug-in SPECTRALMODEL to fit the red noise model (power law model of red noise: $P(f)={A}{[1+(f/f_{\rm c})^{2}]^{(-\alpha/2)}}$, where $A$, $\alpha$, and $f_{\rm c}$ are amplitude, spectral index, and corner frequency, respectively), then referred to this model to obtain the accurate position. After completing the above steps, we employed TEMPO2 to fit ToAs to obtain the best rotation phase model of pulsars. By subtracting the actual observed ToAs from the ToAs predicted by the model, timing residuals were obtained.
@@
Timing residuals were obtained by comparing observations with model-predicted ToAs. We can use TEMPO2 to visualize timing residuals. In an ideal state, the timing residuals should be randomly distributed around zero. If, upon visual inspection, the residuals appear discontinuous or show significant deflection, it may indicate the occurrence of a glitch event in the pulsar. Assuming a glitch event occurs at epoch $t_{\rm g}$, it may result in a change in the pulse phase is $\Delta\phi$, a permanent increase in the spin frequency, $\Delta\nu_{\rm p}$, and a permanent increase in its first derivative $\Delta\dot{\nu}_{\rm p}$. It should be noted that the post-glitch recovery process typically involves one or more exponential decay processes. If the frequency change during the exponential recovery process of the $i$-th phase is $\Delta\nu_{\rm d_{i}}$, and the corresponding timescale is $\tau_{\rm d_{i}}$, we can describe the change of post-glitch pulse phase according to the above-mentioned relevant glitch parameters as (Edwards et al., 2006; Yu et al., 2013):
@@
Table 4: Vortex creep model fit results for 9 large glitches in three gamma-ray pulsars analysed in this work: The characteristic relaxation time of exponential decay $\tau_{\rm creep,exp}$ from the linear response regime in vortex creep model can be confronted with that of the TEMPO2 fitting $\tau_{\rm d}$ listed in Table 3. $\tau_{\rm nl}$ ($\tau_{\rm s}$ in the case of single non-linear regime) is the non-linear creep relaxation time. Post-glitch recovery is expected to end when the relaxations of the non-linear creep regions are over, and glitch induced changes in the spin-down rate recover completely after a waiting time $t_{0}$. Thus the theoretical expectation $t_{\rm th}=\mbox{max}(t_{0},t_{\rm single})$, i.e., the maximum of Eq. (11) and Eq. (13) gives a prediction for the inter-glitch time.
@@ 3 Glitch trigger and post-glitch relaxation in the vortex creep model
The observations of pulsar glitches provided the first evidence of superfluidity in neutron-star interiors (Baym et al., 1969). Previously, our works focused on the amplitude of the glitch and explored whether modern equation of state (EOS) of pulsars can reconcile with the available data (Li, 2015; Li et al., 2016), mainly in the framework of the two-component model (Baym et al., 1969; Anderson & Itoh, 1975), i.e., the superfluid component and the other is the charged component. With various post-glitch recovery data available (Ge et al., 2020; Gügercinoğlu et al., 2022; Liu et al., 2024; Tuo et al., 2024), we here adopt the vortex creep model and confront it with both the observed amplitudes and post-glitch decays (Shang & Li, 2021), in our endeavor to search for a firmly grounded paradigm of glitch phenomenon.
@@
Any observed value of $|\dot{\Omega}_{\mathrm{c}}|$ can represent $|\dot{\Omega}|_{\infty}$ after considering the small magnitude $\Delta\dot{\Omega}_{\mathrm{c}}/\dot{\Omega}_{\mathrm{c}}\sim 10^{-2}$. In the non-linear regime, the interior temperature is sufficiently low compared to the pinning energy, hence we have $\omega_{\infty}\lesssim\omega_{\mathrm{cr}}$ and $\delta\omega_{\rm k}(0)\approx\delta\Omega_{\rm s,k}$. Eq. (10) has the form of making a “shoulder” such that at about $t_{\mathrm{0,k}}$ after the glitch, the spin-down rate decreases rapidly by $I_{\rm s,k}/I$ in a time $\sim 3\tau_{\mathrm{n,k}}$ quasi exponentially. This response resembles the behaviour of the Fermi distribution function of statistical mechanics for non-zero temperatures (Alpar et al., 1984). Single superfluid region’s recovery is thus completed on a timescale,
@@
where $\tau_{\rm sd}=\nu/(2\dot{\nu})$ is the spin-down timescale (characteristic age) of the pulsar. In the limit of $t_{\mathrm{0,A}}\gg\tau_{\mathrm{n,A}}$, Eq. (14) reduces to
@@
with a constant, anomalously large $\ddot{\Omega}_{\mathrm{c}}=(I_{\rm A}/I)(|\dot{\Omega}_{\rm c}|/t_{\rm 0,A})$ (Alpar & Baykal, 2006). The post-glitch recovery is expected to come to end when the relaxations of the non-linear creep regions are over. So the theoretical expectation $t_{\rm th}=\mbox{max}(t_{\mathrm{0,A}},t_{\rm single})$, i.e. maximum of Eq. (11) and Eq. (13) gives a prediction for the inter-glitch time. In practice, we usually choose a summation of several post-glitch recoveries of linear and non-linear regions to fit the post-glitch observations, e.g. a linear region with $I_{\rm{exp}}$ and $\tau_{\rm{creep,exp}}$, a single non-linear region with $I_{\rm{s}}$, $t_{\rm{s}}$ and $\tau_{\rm{s}}$, and a continuous nonlinear regions with $I_{\rm{A}}$, $t_{0}$, and $\tau_{\rm{nl}}$. $\Delta\dot{\Omega}_{\mathrm{c}}/\dot{\Omega}_{\mathrm{c}}$ is related to $\Delta\dot{\nu}/\dot{\nu}$ by $\Omega=2\pi\nu$.

### v2

@@ Abstract
Pulsar glitch is a phenomenon characterized by abrupt changes in the spin period over less than a minute. We present a comprehensive analysis of glitches in four gamma-ray pulsars by combining the timing observation data from Fermi Large Area Telescope (Fermi-LAT) and Parkes 64 m radio telescope. The timing data of five pulsars, namely PSRs J1028--5819, J1420--6048, J1509--5850, J1709--4429 (B1706--44) and J1718--3825, are examined over 14 yr of observations for each. A total of 12 glitches are identified in four pulsars, including a previously unreported glitch. That is, a new small glitch is identified for PSR J1718--3825 in MJD $\sim$ 59121(8), with a fractional glitch size of $\Delta\nu/\nu\sim 1.9(2)\times 10^{-9}$. For PSR J1420--6048, our investigation confirms the presence of two linear recovery terms during the evolution of $\dot{\nu}$ following glitches 4, 6 and 8. Moreover, an exponential recovery process was identified after glitch 8, with a recovery fraction ($Q$) of $Q=0.0131(5)$ and a corresponding timescale of $\tau_{\rm d}=100(6)$ d. Regarding the fourth glitch of PSR J1709--4429, our analysis reveals the presence of two exponential recovery terms with degree of recovery and decay time-scales $Q$1 = 0.0104(5), $\tau_{\rm d1}=72(4)$ d and $Q$2 = 0.006(1), $\tau_{\rm d2}=4.2(6)$ d, respectively. For the remaining previously reported glitches, we also refine the glitch epochs and recovery process through precise fitting of the timing data. We discuss how multi-band data of glitches can help better characterize the glitch recoveries and constrain the underlying physics of glitch events. Our findings demonstrate that the accumulation of observational data reveals the rich complexity of the glitch phenomenon, aiding in the search for a well-established interpretation.
@@
PSR RA DEC $P$ $\dot{P}$ $B_{\rm s}$ $\tau_{\rm c}$ DM Fermi-LAT span Parkes span
@@
1 Introduction
[…]
Young gamma-ray pulsars, characterised by showing high spin-down rates active in the gamma-ray band of emission, are good target sources for searching and studying glitches (Sokolova & Panin, 2022). More than 150 young gamma-ray pulsars have been reported so far, with more than 80 among them detected also in radio emissions (Smith et al., 2023). Investigating glitches in gamma-ray pulsars holds the potential to not only provide valuable insights into the internal structure of neutron stars but also to offer a window into the dynamic evolution of the magnetosphere, see e.g., Kou et al. (2018).
Currently, there has been notable progress in research focused on glitches in gamma-ray pulsars. PSR J2021+4026 is the only glitching pulsar in which a gamma-ray flux jump has been observed (Allafort et al., 2013; Takata et al., 2020). The studies by Ng et al. (2016) and Zhao et al. (2017) indicated that the gamma-ray flux jumps and the corresponding substantial shifts in the pulse profile of PSR J2021+4026 might have a plausible association with glitches. In a similar manner, Lin et al. (2021) reported the detection of a new glitch in PSR J1420--6048 using Fermi-LAT 11 yr timing data and found that the gamma-ray pulse profile changed significantly before and after glitches 4 and 8. Gügercinoğlu et al. (2022) identified 20 glitches in 4 gamma-ray pulsars, and made a relatively accurate prediction of the waiting time for the next glitch through the vortex creep model.
In this paper, we study the glitches of five gamma-ray pulsars by combining the timing observation data of the Fermi-LAT and Parkes telescopes over the past 14 yr. We have identified 12 glitches in total and detected a new small glitch in PSR J1718--3825. Besides, we update the epoch and other parameters of the previously reported 11 glitches.
The structure of this paper is as follows: In Section 2, we briefly introduce the timing observations from Fermi-LAT and Parkes 64 m radio telescope and give an overview of the glitch data processing procedure. One superfluid model, namely the vortex creep model, that is used in glitch fitting is introduced in Section 3, following which we relate the glitch observables to the dynamical and structural properties of the neutron star. The results and underlying physical process of pulsar glitches studied in the paper are detailed in Section 4. We then discuss our results more extensively in Section 5 and finally summarize our findings in Section 6.
@@ 2 Observation and analysis
In this study, we focus on the multi-band research of pulsar glitches. The selected pulsars for this study must meet two criteria simultaneously: first, they exhibit emissions in both radio and gamma-ray wavelengths; second, they have more than 10 yr of high-quality timing data available through long-term monitoring by the Fermi-LAT and the Parkes 64 m radio telescope. Based on these criteria, we selected five young gamma-ray pulsars with spin evolution similar to Vela pulsar for this study, namely PSRs J1028--5819, J1420--6048, J1509--5850, J1709--4429 and J1718--3825, with timing data spanning from 2007 to 2022. Table 1 presents the detailed parameters of pulsars. The first three columns are the pulsar name (J2000) and the position [right ascension (RA) and declination (DEC)] in the equatorial coordinate system. The fourth and fifth columns are the pulsar spin period ($P$) and its first time derivative ($\dot{P}$). The sixth and seventh columns are the inferred surface dipole magnetic field strength at the stellar equator ($B_{\rm s}=3.2\times 10^{19}\sqrt{P\dot{P}}$ G) and the characteristic age ($\tau_{\rm c}=P/2\dot{P}$) of pulsars. The last three columns are the pulsar’s dispersion measure (DM) and the timing observation data time span of the Fermi-LAT and the Parkes telescopes for the corresponding sources.
@@
We also collected the timing observation data of these five gamma-ray pulsars from 2007 to 2020 on the Parkes pulsar data archive [^2] https://data.csiro.au/domain/atnf. Manchester et al. (2013) has introduced the timing observation of Parkes 64 m radio telescope in detail. Briefly, Parkes observation was carried out at 20 cm band, mainly using the multi-beam receiver, with a centre frequency of 1369 MHz and bandwidth of 256 MHz (Staveley-Smith et al., 1996; Hobbs et al., 2020). The Parkes Digital Filter Bank systems (including PDFB1, PDFB2, PDFB3, and PDFB4) are used for data recording. The raw data consists of 1024 frequency channels, each with a bandwidth of 0.25 MHz. The integration times were 2 to 15 mins, with a sub-integration time of 30 s and the observation intervals of 2 to 4 weeks for each pulsar.
@@
With regarding to the radio observations, the PSRCHIVE package was used to process the raw data. After eliminating radio frequency interference (RFI), data are de-dispersed in frequency and summed in frequency and time to obtain an integrated pulse profile for each observing session (Hotan et al., 2004; van Straten et al., 2012). The PAT tool of PSRCHIVE was then used to cross-correlate each mean profile with the standard profile, leading to produce the time-of-arrivals (ToAs; Taylor 1992). The standard profile is formed by aligning the pulse phase and summing all the observations. Next, the Jet Propulsion Laboratory planetary ephemeris DE440 (Park et al., 2021) and the Barycentric Coordinate Time (TCB) were used to convert these local ToAs to (Solar System Barycenter) SSB, thereby eliminating the effect of the Earth motion on the ToAs.
@@
When combining the gamma-ray and radio ToAs, to ensure the phase alignment of the data from two telescopes, the package TEMPO2 (Hobbs et al., 2006; Edwards et al., 2006) has been used to fit the phase offset of the data of the Parkes telescope and Fermi-LAT. The phase offset value obtained from this fitting will be appended to the ephemeris to achieve phase alignment of the timing data of the two telescopes. We also introduced two parameters, EFAC and EQUAD, to correct the uncertainty of each ToA acquired from different telescopes (Wang et al., 2015). Among them, EFAC is a free parameter that adjusts unaccounted instrumental errors, and EQUAD accounts for additional time-independent white noise processes (Lentati et al., 2014). The values of these two parameters were determined by the EFACEQUAD plug-in of TEMPO2, and this method did not involve Bayesian techniques. After correction, the relationship between the initial uncertainty ($\delta$) of ToAs and its new value (${\delta}_{\rm s}$) was defined as ${\delta}_{\rm s}^{\rm 2}={\rm EFAC}^{2}\times(\delta^{2}+{\rm EQUAD}^{2})$. After adding the error correction of ToA, the reduced chi-square values of each fitting region are distributed between 0.95–1.05.
The pulsar position is one important factor that probably affects the precision of the glitch parameters, but for some sources, the position obtained from Pulsar Catalogue (PSRCAT; Manchester et al. 2005) [^3] https://www.atnf.csiro.au/research/pulsar/psrcat/ is not accurate enough. Although the accurate positions of most bright pulsars can be determined by VLBI (Very Long Baseline Interferometry), the positions of most relatively weak pulsars could be constrained by pulsar timing. For these pulsars, we used the "Cholesky" method to minimize the influence of timing noise on the determination of the pulsar position (Coles et al., 2011). To do this, we selected a sufficiently long length of data that not include any glitch, invoked the TEMPO2 plug-in SPECTRALMODEL to fit the red noise model (power law model of red noise: $P(f)={A}{[1+(f/f_{\rm c})^{2}]^{(-\alpha/2)}}$, where $A$, $\alpha$, and $f_{\rm c}$ are amplitude, spectral index, and corner frequency, respectively). Then, by analyzing this red noise model and applying a global least-squares fitting procedure, we achieved precise fitting of the pulsar position parameters. For example, the position parameters of PSR J1028--5832 has been re-determined, updating the RA to 10:28:27.9(1) and the DEC to --58:19:06.21(6). These updated values are included in Table 1. Compared to the cataloged position [RA 10:28:28.0(1) and DEC --58:19:05.2(15)] provided by PSRCAT, the RA remains nearly unchanged, while the DEC shows a significant shift of approximately $1^{\arcsec}$. Additionally, the uncertainty in the DEC has been substantially reduced. For the other four pulsars, as their new positions are consistent with previously published data, we used the previously reported positions in the timing analysis. Note that the red noise model is exclusively utilized for deriving the pulsar position parameters and is not applied to other data analysis processes.
After completing the above steps, we employed TEMPO2 to fit ToAs to obtain the best rotation phase model of pulsars. By subtracting the actual observed ToAs from the ToAs predicted by the model, timing residuals were obtained.
@@
We can use TEMPO2 to visualize timing residuals. If the timing model is correct, the timing residuals should be randomly distributed around zero. If, upon visual inspection, the residuals appear discontinuous or show significant deflection, it may indicate the occurrence of a glitch event in the pulsar.
Assuming a glitch event occurs at epoch $t_{\rm g}$, it may result in a change in the pulse phase is $\Delta\phi$, a permanent increase in the spin frequency, $\Delta\nu_{\rm p}$, and a permanent increase in its first derivative $\Delta\dot{\nu}_{\rm p}$. It should be noted that the post-glitch recovery process typically involves one or more exponential decay processes. If the frequency change during the exponential recovery process of the $i$-th phase is $\Delta\nu_{\rm d_{i}}$, and the corresponding timescale is $\tau_{\rm d_{i}}$, we can describe the change of post-glitch pulse phase according to the above-mentioned relevant glitch parameters as (Edwards et al., 2006; Yu et al., 2013):  [^4] Although $\ddot{\nu}$ may vary following a glitch, this variation is likely contaminated by the post-glitch recovery process, $\Delta\ddot{\nu}$ was excluded from Eq. (2).
@@
Table 4: Vortex creep model fit results for 9 large glitches in three gamma-ray pulsars analysed in this work: The characteristic relaxation time of exponential decay $\tau_{\rm creep,exp}$ from the linear response regime in vortex creep model can be confronted with that of the TEMPO2 fitting $\tau_{\rm d}$ listed in Table 3. $\tau_{\rm nl}$ ($\tau_{\rm s}$ in the case of single non-linear regime) is the non-linear creep relaxation time. Post-glitch recovery is expected to end when the relaxations of the non-linear creep regions are over, and glitch induced changes in the spin-down rate recover completely after a waiting time $t_{0}$. Thus the theoretical expectation $t_{\rm th}=\mbox{max}(t_{0},t_{\rm single})$, i.e., the maximum of Eq. (11) and Eq. (13) gives a prediction for the inter-glitch time. The observed inter-glitch time $t_{\mathrm{obs}}$ is given for comparison with the theoretical prediction by the vortex creep model $t_{\rm th}$.
@@ 3 Glitch trigger and post-glitch relaxation in the vortex creep model
The observations of pulsar glitches provided the first evidence of superfluidity in neutron-star interiors (Baym et al., 1969). Previously, our works focused on the amplitude of the glitch and explored whether modern equation of state (EOS) of pulsars can reconcile with the available data (Li, 2015; Li et al., 2016), mainly in the framework of the two-component model (Baym et al., 1969; Anderson & Itoh, 1975), i.e., the (crustal) superfluid component and the other is the charged component consisting of all matter in the normal phase and the proton-neutron fluid conglomerate coupled to each other via electrons scattering off vortex lines and strong magnetic field. With various post-glitch recovery data available (Ge et al., 2020; Gügercinoğlu et al., 2022; Liu et al., 2024), we here adopt the vortex creep model and confront it with both the observed amplitudes and post-glitch decays (Shang & Li, 2021), in our endeavor to search for a firmly grounded paradigm of glitch phenomenon.
@@
Because $\dot{\Omega}_{\rm s}=\dot{\Omega}_{\rm c}=\dot{\Omega}_{\infty}$ in the steady state and the magnitude of $\Delta\dot{\Omega}_{\rm c}$ is small compared to $\dot{\Omega}_{\rm c}$, $\Delta\dot{\Omega}_{\mathrm{c}}/\dot{\Omega}_{\mathrm{c}}\sim 10^{-2}$, we use observed value of $|\dot{\Omega}_{\rm c}|$ to represent $|\dot{\Omega}|_{\infty}$. In the non-linear regime, the interior temperature is sufficiently low compared to the pinning energy, hence we have $\omega_{\infty}\lesssim\omega_{\mathrm{cr}}$ and $\delta\omega_{\rm k}(0)\approx\delta\Omega_{\rm s,k}$. Eq. (10) has the form of making a “shoulder” such that at about $t_{\mathrm{0,k}}$ after the glitch, the spin-down rate decreases rapidly by $I_{\rm s,k}/I$ in a time $\sim 3\tau_{\mathrm{n,k}}$ quasi exponentially. This response resembles the behaviour of the Fermi distribution function of statistical mechanics for non-zero temperatures (Alpar et al., 1984). Single superfluid region’s recovery is thus completed on a timescale,
@@
where $\tau_{\rm c}=\nu/(2\dot{\nu})$ is the spin-down timescale (characteristic age) of the pulsar. In the limit of $t_{\mathrm{0,A}}\gg\tau_{\mathrm{n,A}}$, Eq. (14) reduces to
@@
with a constant, anomalously large $\ddot{\Omega}_{\mathrm{c}}=(I_{\rm A}/I)(|\dot{\Omega}_{\rm c}|/t_{\rm 0,A})$ (Alpar & Baykal, 2006). The post-glitch recovery is expected to come to end when the relaxations of the non-linear creep regions are over. So the theoretical expectation $t_{\rm th}=\mbox{max}(t_{\mathrm{0,A}},t_{\rm single})$, i.e. maximum of Eq. (11) and Eq. (13) gives a prediction for the inter-glitch time. In practice, we usually choose a summation of several post-glitch recoveries of linear and non-linear regions to fit the post-glitch observations, e.g., a linear region with $I_{\rm{exp}}=I_{1}$ and $\tau_{\rm{creep,exp}}=\tau_{\rm l,1}$ in Eq. (7), a single non-linear region with $I_{\rm{s}}=I_{\rm s,1}$, $t_{\rm{s}}=t_{0,1}$ and $\tau_{\rm{s}}=\tau_{\rm n,1}$ in Eq. (10), and a continuous nonlinear regions with $I_{\rm{A}}$, $t_{0}=t_{0,\rm A}$, and $\tau_{\rm{nl}}=\tau_{\rm{n,A}}$ in Eq. (14). $\Delta\dot{\Omega}_{\mathrm{c}}/\dot{\Omega}_{\mathrm{c}}$ is related to $\Delta\dot{\nu}/\dot{\nu}$ by $\Omega=2\pi\nu$.